Please help -
Alicia can row 7 miles downstream in the same time it takes her to row 4 miles upstream. She rows downstream 5 miles/hour faster than she rows upstream. Find Alicia’s rowing rate each way.

Answer by ptaylor(2198)

Distance(d) equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
Let r=Alicia's rate upstream
Then r+5=her rate downstream
Time to row 4 miles upstream=4/r
Time to row 7 miles downstream=7/(r+5)
And we are told that these times are the same, sooooo
4/r=7/(r+5) multiply each term by r(r+5) or simply cross multiply
4(r+5)=7r
4r+20=7r
3r=20
r=6 2/3 hr-----------rowing rate upstream
r+5=6 2/3 +5=11 2/3 hr---rowing rate downstream
CK
t=d/r
time upstream=4/(6 2/3)=4/(20/3)---complex fraction; make it a simple fraction by multiplying the numerator and denominator by (3/20) and we get (4*(3/20))/(20/3*3/20)=12/20=3/5 hr
time downstream=7/(11 2/3)=7/(35/3);(7*(3/35))/(35/3*3/35)=21/35=3/5 hr
OK
Hope this helps---ptaylor